If 3x+4y= -2 and x-6y=-8, then the value of x is?
a) -1
b) 1
c) 2
d) -2
e) 0
I would appreciate your help, thanks :)

Answer by mananth(16946)

3 x + 4 y = -2 .............1
1 x -6 y = -8 .............2

multiply (1)by 6
Multiply (2) by 4
18 x + 24 y = -12
4 x -24 y = -32
Add the two equations
22 x = -44
/ 22
x = -2
plug value of x in (1)
3 x + 4 y = -2
-6 + 4 y = -2
4 y = -2 + 6
4 y = 4
y = 1